OpenStudy (anonymous):

Factor the following problem completely. First, factor out the greatest common factor, and then factor the remaining trinomial: 2y^4 − 4y^3 − 30y^2

OpenStudy (cwrw238):

GCF = 2y^2

OpenStudy (cwrw238):

2y^2(y^2 - 2y - 15) now you need to factor the trinomial in the parentheses

OpenStudy (anonymous):

I have no idea what to do next...

OpenStudy (cwrw238):

ok tghis will factor to the form (x + a)(x + b) where ab = the last term ( -15 ) and a + b = -2 ( the coefficient of 2) ab = -15 one possibility = -5 and +3 what is the value of a + b ?

OpenStudy (anonymous):

2y^2(y^2 - 2y - 15)

OpenStudy (cwrw238):

a + b = -5 + 3 = -2 so that it factors are 2y^2(y - 5)(y + 3)

OpenStudy (anonymous):

oooooh ok, thank you !

OpenStudy (cwrw238):

yw to get the factors you do this -5 and + 3 give us the 2 middle terms -5y + 3y y^2 - 2y - 15 = y^2 -5y + 3y - 15 grouping in pairs: = y(y - 5) + 3 (y - 5) = (y + 3)(y - 5)

OpenStudy (anonymous):

So, if you add it together, it equals the middle term, but when you multiply it, it equals the last term?

OpenStudy (cwrw238):

right

OpenStudy (cwrw238):

eg x^2 - 5x + 6 what do you think the 2 numbers would be in this case ?